OpenStudy (anonymous):

what is the equation of the line that passes through the points -1,3/2 and 1/2,1/2

OpenStudy (tester97):

Well since we know that y intercept form is Y=mx+b can you plug in the variables here?

OpenStudy (anonymous):

need help

OpenStudy (tester97):

I will not give you the answer without you putting some effort into trying. Just give it a go. You can do it if you put your mind to it!

OpenStudy (texaschic101):

we need to find the slope first by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-1, 3/2)....x1 = -1 and y1 = 3/2 (1/2, 1/2)...x2 = 1/2 and y2 = 1/2 now sub slope(m) = (1/2 - 3/2) / (1/2 - (-1) slope(m) = (1/2 - 3/2) / (1/2 + 1) can you find the slope now ? Once you give it a try and let me see what you can do, then I will continue to help

OpenStudy (tester97):

Thank you @texaschic101 but this person just want flat out answers. http://prntscr.com/42k8h1 as you can see they posted another question

OpenStudy (texaschic101):

not gonna happen ...at least not from me

OpenStudy (anonymous):

nooo ii dont soo please mind your bussiness ii have the answer so thankss

OpenStudy (tester97):

You just proved my point......

OpenStudy (anonymous):

yesss soo get it correct bye

OpenStudy (tester97):

wow rude!